A year ago, a mother was eight as old as her daughter. Now, her age is the square of her daughter's age. How old are they now?

Answer by scott8148(6628)

"A year ago, a mother was eight as old as her daughter" ___ m - 1 = 8(d - 1) ___ m = 8d - 7

"her age is the square of her daughter's age" ___ m = d^2

substituting ___ (8d - 7) = d^2 ___ 0 = d^2 - 8d + 7

factoring ___ 0 = (d - 7)(d - 1)